OpenStudy (anonymous):

What is the value of the expression? 7^-3 x 7^9 over 7^-8

imqwerty (imqwerty):

use these identities-\[a^x \times a^y=a^{x+y}\]and this one too-\[\frac{ a^x }{ a^y }=a^{x-y}\]

OpenStudy (anonymous):

a. 7^14 b. 7^2 c. 7^-14 d. 7^-2

imqwerty (imqwerty):

did u try it?

OpenStudy (anonymous):

so its c? @imqwerty

imqwerty (imqwerty):

no :) check ur calculations once again

OpenStudy (anonymous):

a?

imqwerty (imqwerty):

yess!! :D

OpenStudy (pawanyadav):

8^(-3+9+8)=8^14 Can you tell why +8?